Smith & Sons has a debt-equity ratio of.55. What is the total debt ratio? Select one: a. .46 b. .55 c. .51 d. .49 e. .35

Glen Acre Wines has sales of $682,100, total debt of $285,000, total equity of $323,900, and a profit margin of 8 %. What is the return on assets?

Select one: a. 9.17 % b. 9.11 % c. 6.28 % d. 9.03 % e. 8.96 %
Answer #1

Total Debt ratio = Debt/(Debt+Equity)

= 0.55/(0.55+1)

= 0.35

i.e. e

Return on Assets = Sales*Profit Margin/Total Assets

= 682100*8%/(285000+323900)

= 8.96%

i.e. e.